OpenStudy (maheshmeghwal9):

If sink, cosk are the roots of the equation ax^2+bx+c=0. Then, find a polynomial satisfying the above conditions in terms of a,b,c.

OpenStudy (anonymous):

\[ ax^2+bx=c=0 \] Please check again.

OpenStudy (anonymous):

is it - or + ?

OpenStudy (maheshmeghwal9):

+

OpenStudy (maheshmeghwal9):

sorry

OpenStudy (anonymous):

a(x-cosk)(x-sink)=0

OpenStudy (maheshmeghwal9):

now what am i gonna do?

OpenStudy (anonymous):

That's your polynomial. simplify it now :)

OpenStudy (maheshmeghwal9):

Answer is \[(a+c)^2=b^2+c^2.\]

OpenStudy (anonymous):

Sorry It's polynomial, so it should be just a(x-cosk)(x-sink)

OpenStudy (maheshmeghwal9):

But I have given answer which is true.

OpenStudy (anonymous):

Got it!

OpenStudy (maheshmeghwal9):

But answer is